OpenStudy (anonymous):

how can i resolve this problem. Two U.S. states are not part of the continental United States. What percent of the fifty U.S. states are included in the continental United States?

OpenStudy (geneticrockhopper247):

4% are part of. 96% are part of. 50*2 is 100 and all percents add up to 100. So, if there are two states, then two of the states *2 (in order to get the percent) is 4%. There are 48 other states, so 48*2=96%. Does that make sense???

OpenStudy (anonymous):

yes thank you

OpenStudy (geneticrockhopper247):

no problem

OpenStudy (texaschic101):

There is 50 U.S. states. Thats 100 %. 48 states are part of the continental U.S. so 48/50 = 24/25 = .96 = 96% 2 states are not part of the U.S., so 2/50 = 1/25 = .04 = 4%